Suppose A and B are n x n matrices with A invertible. Prove that det ABA^-1 = det B
1
Expert's answer
2020-05-11T18:02:40-0400

matrices are not commutative. that is AB is not equal to BA

"det(BA)=det(B)det(A)\\det(BA)=det(A)det(B)\\det(BA)=det(AB)"

"BB^{-1}=I \implies det(I)=1"


hence to show "det (ABA^{-1}) = det( B)"

"\implies det (ABA^{-1}) =det(A)det(B)det(A^{-1})\\det (ABA^{-1}) =det(A)det(A^{-1})det(B)\\
det (ABA^{-1})=(det(A)det(A^{-1}))det(B)\\det (ABA^{-1})=(det(AA^{-1}))det(B)"


but "AA^{-1}=I"

"det (ABA^{-1})=(det(I))det(B)\\" "det(I)=1"

hence:


"det (ABA^{-1})=det(B)"